OpenStudy (anonymous):

PLZ HELP?!?!?!?!(WILL MEDAL+FAN) The functions p(x) and g(x) are shown below: g(x) = 0.09x p(x) = (0.09)x Which statement best describes the graph of p(x) and g(x)? A. The graphs will both have their y-intercept equal to 1. B. The graphs will both have their y-intercept equal to 0.09. C. The graph of p(x) will eventually exceed the graph of g(x). D.The graph of g(x) will eventually exceed the graph of p(x).

OpenStudy (anonymous):

i dont know how to to this..

OpenStudy (amilapsn):

Did you mean this?\[g(x)=0.09x\\ p(x)=0.09^x\]

OpenStudy (anonymous):

it says g(x) = 0.09x p(x) = (0.09)x

OpenStudy (anonymous):

i think its either a or b

OpenStudy (freckles):

ab is the same as a(b) I'm kind of thinking you meant what @amilapsn wrote above

OpenStudy (anonymous):

yeah they do mean the same thing

OpenStudy (freckles):

\[\\ p(x)=0.09^x \text{ this is the same as } \\ p(x)=(0.09)^x \\ \text{ if you have happen to have ( )} \\ \text{ in your problem and you don't know what they mean }\] so did @amilapsn interpret your problem correctly

OpenStudy (anonymous):

yes, i agree it can be written p(x)=0.09^x

OpenStudy (freckles):

ok so great we are in agreement (0.09)x is not the same as (0.09)^x correct?

OpenStudy (freckles):

anyways do you know how to graph the line and the exponential function ?

OpenStudy (anonymous):

no

OpenStudy (freckles):

A. The graphs will both have their y-intercept equal to 1. can we answer this question then? to find the y-intercept set x=0 and find y. \[\\ \text{ for } g \text{ we have } y=0.09(0)=? \\ \text{ for } p \text{ we have } y=0.09^0=?\] are the y-intercepts the same?

OpenStudy (anonymous):

Yes they both equal 1 oh i was over thinking it all i had to do was solve for y Thank you, ur a life saver

OpenStudy (freckles):

no

OpenStudy (freckles):

0.09 times 0 is 0

OpenStudy (freckles):

0.09 to the 0 power is 1

OpenStudy (freckles):

they are not both the same so choice A is definitely not the answer let's look at choice B

OpenStudy (freckles):

B. The graphs will both have their y-intercept equal to 0.09. we know this is false too because we already found the y-intercepts previously

OpenStudy (freckles):

so the next thing to do is see which function eventually exceeds the other

OpenStudy (freckles):

and it would be kinda helpful to graph/draw the functions

OpenStudy (freckles):